The sum of three consecutive number is 9. Find the numbers
Answer by ikleyn(52781)

.

(n-1) + n + (n+1) = 9


       3n         = 9

        n         = 9/3 = 3


The numbers are  2,3,4.       ANSWER

Solved, answered and explained.


            Notice and MEMORIZE:


the sum of any three consecutive integer numbers is  ALWAYS  the tripled mid number.